The most recent financial statements for Bello Co. are shown here: Income Statement Sales Costs Taxable Balance Sheet $11,700 Deb15,700 12,800 Fixed assets 26,500 Equity 22,500 18,900 Current assets 6,100 38,200 Tota $38,200 income Taxes (21%) 1,281 Net income$ 4,819 Assets and costs are proportional to sales. Debt and equity are not. The company maintains a constant 30 percent dividend payout ratio. What is the sustainable growth rate? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) Sustainable growth rate
Answer #1

Answer:

Sustainable Growth rate = ROE * b / (1 – ROE * b)
ROE = Net Income / Equity * 100
ROE = $4,819 / $22,500 * 100
ROE = 21.4178%

Retention Ratio (b) = 1 – Dividend Payout Ratio
Retention Ratio (b) = 1 – 0.30
Retention Ratio (b) = 0.70 or 70%

Sustainable Growth rate = 0.214178 * 0.70 / (1 – 0.214178 * 0.70)
Sustainable Growth rate = 0.1499246 / (1 – 0.1499246)
Sustainable Growth rate = 0.1499246 / 0.8500754
Sustainable Growth rate = 17.64%